Suppose that the economy is characteised by the following behavioral equations C=Co +c1Yd
Where : Co=280
I=270
G=300
T=200
Marginal propensity to save (MPS) is 0.4(or 40%)
Solve for
a) equlibrum Gdp (Y)
b) disposable income (Yd)
C) Consumption spending (C)
1
Expert's answer
2020-04-23T11:51:40-0400

"a) c1=1-MPS; c1=1-0.4=0.6;\nYd=Y-T;"

"Y=C+I+G;\nY=280+0.6\u00d7(Y-200)+270+300; Y=730+0.6Y;"

"0.4Y=730;"

"Y=" $ "1825;"

"b)Yd=1825-200=" $"1625;"

"c) C=280+0.6\u00d71625=" $ "1255;"